OpenStudy (anonymous):

Simplify quantity 6 x plus 9 over 15 x squared all over quantity 16 x minus 12 over 10 x to the fourth power

OpenStudy (anonymous):

Sooo \[\large \sf \frac{6x+\frac{9}{15}x^{2}}{16x-\frac{12}{10}x^{4}}\] is it?

OpenStudy (anonymous):

Remember that a complex fraction can be rewritten as the top fraction multiplied by the reciprocal of the bottom fraction: \[\frac{ \frac{ a }{ b } }{ \frac{ c }{ d } }=(\frac{ a }{ b })(\frac{ d }{ c })\] Try to do that with this fraction and it should get much easier to cancel things.

jhonyy9 (jhonyy9):

- using what above @Alphabet_Sam have wrote,will get 6x +9 -------- 15x^2 6x+9 10x^4 ---------- = --------- * --------- = ? 16x-12 15x^2 16x-12 --------- 10x^4 - do you can continue it ?